# Nested Sampling for Exploring Lennard - Jones Clusters

### v1

@@ 4 Program profiling and optimisation
In this section, we look at the impact of the different developments made to nested_fit on the computational resources needed. For that, we take the 36-atoms cluster for its high number of parameters (108). First, we look at the result of profiling — performed with valgrind and gprof2dot —, to see which parts of the algorithm are the most computationally expensive. We used $K=500$ and $n_{bases}=1$. We consider three cases: slice sampling transformed with the calculation of the covariance matrix at each iteration, slice sampling transformed with the calculation of the covariance matrix every $0.05K$ iterations and slice sampling with the calculation of the covariance matrix every $0.05K$ iterations. The results of the profiling are presented in Figures 3, 4 and 5 respectively. In the first case, version 4.5.5 of nested_fit was used and version 4.6.0 for the two other cases.
@@
The purpose of each function is explained in Table 1. Only the relevant boxes are represented. The last three lines in a box corresponds to (1) the percentage of time spent in this function and its children function, (2) the percentage of time spent in this function alone and (3) the number of times the function was called. The arrows linking one box to another represent the first (parent) function calling the second (child) with the two numbers being the percentage of time the child function transfers to its parent and the number of times the parent function calls the child function. A box is only represented if at least 5% of the time is spent in the function.
@@
First, looking at Figures 3 and 4, we can see that, in the case of slice sampling transformed, over 50% of the time is spent going back to the real space to calculate the likelihood and check the boundaries (i.e. in the part_like_sub and test_bnd_sub functions) which is not needed in slice sampling. Thus, more time is spent changing space than calculating the function studied, leading to a computational overhead for slice sampling transformed. Furthermore, we can see that when the covariance function is computed at each iteration (Figure 3), this calculation takes around 20% of the time while when it is only computed every $0.05K$ iterations (here every 25 iterations) (Figures 4 and 5), the box does not appear, which means that less than 5% of the time is spent on this particular task. Finally, we see that, for slice sampling transformed with the covariance matrix computed at each iteration (Figure 3), around 12% of the time is spent calculating the function studied. This number goes up to around 55% (4.4 times more) when using slice sampling and computing the covariance function every $0.05K$ iterations (Figure 5). In both cases, the number of times the energy function was called is roughly the same (around 580000 times). This shows that, with our optimisations, we removed calculations that are not needed and that bring a significant computational overhead. This is further shown by Table 2 that gives the computational time of all three cases for 20000 iterations: with the optimisations, the computational time is reduced by a factor 2.8.
@@
Figure 5: Profiling for slice sampling with the calculation of the covariance matrix every $0.05K$ iterations
@@
For the impact of the parallelisation on the computation time, we also look at the time required to run nested_fit with and without parallelisation using slice sampling transformed with $K=1000$ and $n_{bases}=1$. The covariance matrix is computed every iteration. The parallelised case over 64 cores is around 21 times faster than the non-parallelised case (the non-parallelised case takes 85 minutes while the parallelised case takes 4 minutes). A possible explanation for this dividing factor being smaller than the number of cores used is that the time gained by simultaneously searching 64 new points may be partially counterbalanced by the time taken for the sequential adding of all the points, some of which may be rejected due to the constraint changing with each added point.
@@
Performs one iteration of nested sampling with slice sampling
@@
Table 1: Purpose of the functions shown in Figures 3 to 5.
Version 4.5.5 Version 4.6.0 Version 4.6.0
Slice sampling Slice sampling Slice sampling
transformed transformed
@@ 5 Conclusion
First, we have studied clusters of two sizes. As for the 7-atoms cluster, we have seen that we were able to recover phase transitions by comparing our work with that in Ref. [10]. We saw that we required more live points and more likelihood calls that in Ref. [10], likely due to the different choices of method to find a new point, which is an inconvenient for a more computationally expensive energy function. We then studied the 36-atoms cluster which presents a solid-solid phase transition at low temperature that nested_fit was able to recover provided enough live points are used.
We have then looked at the impact of different implementation on the computational resources: we have seen that using slice sampling transformed brought a consequent computational overhead through the need to transform back to the real space to compute the energy. Slice sampling, which works directly in the real space, does not need to perform this task and therefore spends more than triple the time calculating the energy compared to slice sampling transformed, even though the function was called around the same number of times. We also saw that, using parallelisation on 64 cores, we gained a factor 21 in the computation time. All of those developments allowed us to study bigger and more complex clusters that require a higher number of live points. For example, one run for the 36-atoms Lennard - Jones clusters with $K=70000$ took around 42 hours to run with the parallelisation.
@@
Acknowledgments
Work realised with the support of the Sorbonne Center for Artificial Intelligence 579 - Sorbonne University - IDEX SUPER 11-IDEX-0004. The authors thank Philipe Depondt, Simon Huppert and Julien Salomon for helpful discussions.
Références
  • [1] N. W. Ashcroft and N. D. Mermin. Solid state physics. Brooks/Cole, 1976.
  • [2] J. P. K. Doye, M. A. Miller, and D. J. Wales. Evolution of the potential energy surface with size for Lennard-Jones clusters. The Journal of Chemical Physics, 111(18):8417–8428, 1999.
  • [3] D. J. Wales and J. P. K. Doye. Global Optimization by Basin-Hopping and the Lowest Energy Structures of Lennard-Jones Clusters Containing up to 110 Atoms. The Journal of Physical Chemistry A, 101(28):5111–5116, 1997.
  • [4] N. Quirke and P. Sheng. The melting behavior of small clusters of atoms. Chemical Physics Letters, 110(1):63–66, 1984.
  • [5] T. L. Beck and R. S. Berry. The interplay of structure and dynamics in the melting of small clusters. The Journal of Chemical Physics, 88(6):3910–3922, 1988.
  • [6] J. P. Neirotti, F. Calvo, D. L. Freeman, and J. D. Doll. Phase changes in 38-atom Lennard-Jones clusters. I. A parallel tempering study in the canonical ensemble. The Journal of Chemical Physics, 112(23):10340–10349, 2000.
  • [7] F. Calvo, J. P. Neirotti, D. L. Freeman, and J. D. Doll. Phase changes in 38-atom Lennard-Jones clusters. II. A parallel tempering study of equilibrium and dynamic properties in the molecular dynamics and microcanonical ensembles. The Journal of Chemical Physics, 112(23):10350–10357, 2000.
  • [8] J. Skilling. Nested Sampling. In AIP Conference Proceedings, volume 735, pages 395–405, Garching (Germany), 2004. AIP. ISSN: 0094243X.
  • [9] G. Ashton, N. Bernstein, J. Buchner, X. Chen, G. Csányi, A. Fowlie, F. Feroz, M. Griffiths, W. Handley, M. Habeck, E. Higson, M. Hobson, A. Lasenby, D. Parkinson, L. B. Pártay, M. Pitkin, D. Schneider, J. S. Speagle, L. South, J. Veitch, P. Wacker, D. J. Wales, and D. Yallup. Nested sampling for physical scientists. Nature Reviews Methods Primers, 2(1):39, 2022. arXiv:2205.15570.
  • [10] L. B. Pártay, A. P. Bartók, and G. Csányi. Efficient Sampling of Atomic Configurational Spaces. The Journal of Physical Chemistry B, 114(32):10502–10512, 2010.
  • [11] W. Shi and J. K. Johnson. Histogram reweighting and finite-size scaling study of the Lennard–Jones fluids. Fluid Phase Equilibria, 187-188:171–191, 2001.
  • [12] L. B. Pártay, C. Ortner, A. P. Bartók, C. J. Pickard, and G. Csányi. Polytypism in the ground state structure of the Lennard-Jonesium. Physical Chemistry Chemical Physics, 19(29):19369–19376, 2017.
  • [13] L. B. Pártay, G. Csányi, and N. Bernstein. Nested sampling for materials. The European Physical Journal B, 94(8):159, 2021.
  • [14] M. Trassinelli. The Nested_fit Data Analysis Program. In The 39th International Workshop on Bayesian Inference and Maximum Entropy Methods in Science and Engineering, page 14. MDPI, 2019.
  • [15] M. Trassinelli. Bayesian data analysis tools for atomic physics. Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms, 408:301–312, 2017.
  • [16] M. Trassinelli and P. Ciccodicola. Mean Shift Cluster Recognition Method Implementation in the Nested Sampling Algorithm. Entropy, 22(2):185, 2020.
  • [17] L. Maillard, F. Finocchi, and M. Trassinelli. Assessing Search and Unsupervised Clustering Algorithms in Nested Sampling. Entropy, 25(2):347, 2023.
  • [18] R. M. Neal. Slice sampling. The Annals of Statistics, 31(3), 2003.
  • [19] W. J. Handley, M. P. Hobson, and A. N. Lasenby. PolyChord: nested sampling for cosmology. Monthly Notices of the Royal Astronomical Society: Letters, 450(1):L61–L65, 2015. arXiv:1502.01856.
  • [20] W. J. Handley, M. P. Hobson, and A. N. Lasenby. polychord: next-generation nested sampling. Monthly Notices of the Royal Astronomical Society, 453(4):4385–4399, 2015.
  • [21] R. W. Henderson and P. M. Goggans. Parallelized nested sampling. AIP Conference Proceedings, pages 100–105, 2014.
  • [22] A. Hüller. First order phase transitions in the canonical and the microcanonical ensemble. Zeitschrift für Physik B Condensed Matter, 93(3):401–405, 1994.
  • [23] I. Medved’, A. Trník, and L. Vozár. Modeling of heat capacity peaks and enthalpy jumps of phase-change materials used for thermal energy storage. International Journal of Heat and Mass Transfer, 107:123–132, 2017.
  • [24] A. Kokalj. XCrySDen—a new program for displaying crystalline structures and electron densities. Journal of Molecular Graphics and Modelling, 17(3):176–179, 1999.
  • [25] R. J. N. Baldock, L. B. Pártay, A. P. Bartók, M. C. Payne, and G. Csányi. Determining pressure-temperature phase diagrams of materials. Physical Review B, 93(17):174108, 2016. arXiv:1503.03404 [cond-mat, physics:physics].
  • [26] R. J. N. Baldock, N. Bernstein, K. M. Salerno, L. B. Pártay, and G. Csányi. Constant-pressure nested sampling with atomistic dynamics. Physical Review E, 96(4):043311, 2017.
  • [27] V. A. Mandelshtam and P. A. Frantsuzov. Multiple structural transformations in Lennard-Jones clusters: Generic versus size-specific behavior. The Journal of Chemical Physics, 124(20):204511, 2006.
  • [28] P. A. Frantsuzov and V. A. Mandelshtam. Size-temperature phase diagram for small Lennard-Jones clusters. Physical Review E, 72(3):037102, 2005.

### v2

@@ 4 Program profiling and optimisation
In this section, we look at the impact of the different developments made to nested_fit on the computational resources needed. For that, we take the 36-atom cluster for its high number of parameters (108). First, we look at the result of profiling — performed with valgrind noauthor_valgrind_nodate; nethercote_valgrind_2007 and gprof2dot noauthor_gprof2dot_nodate —, to see which parts of the algorithm are the most computationally expensive. We use $K=500$ and $n_{\textrm{bases}}=1$. We consider three cases: slice sampling transformed with the calculation of the covariance matrix at each iteration, slice sampling transformed with the calculation of the covariance matrix every $0.05K$ iterations and slice sampling real with the calculation of the covariance matrix every $0.05K$ iterations. The results of the profiling are presented in Figures 3, 4 and 5 respectively. In the first case, version 4.5.5 of nested_fit is used and version 4.6.0 for the two other cases.
@@
The purpose of each function is explained in Table 1. Only the relevant boxes are represented. The last three lines in a box correspond to (1) the percentage of time spent in this function and its children function, (2) the percentage of time spent in this function alone and (3) the number of times the function was called. The arrows linking one box to another represent the first (parent) function calling the second (child) with the two numbers being the percentage of time the child function transfers to its parent and the number of times the parent function calls the child function. A box is only represented if at least 5% of the time is spent in the function.
@@
First, looking at Figures 3 and 4, we can see that, in the case of slice sampling transformed, over 50% of the time is spent going back to the real space to calculate the likelihood and check the boundaries (i.e. in the part_like_sub and test_bnd_sub functions) which is not needed in slice sampling real. Thus, more time is spent changing space than calculating the function studied, leading to a computational overhead for slice sampling transformed. Furthermore, we can see that when the covariance function is computed at each iteration (Figure 3), this calculation takes around 20% of the time while when it is only computed every $0.05K$ iterations (here every 25 iterations) (Figures 4 and 5), the box does not appear, which means that less than 5% of the time is spent on this particular task. Finally, we see that, for slice sampling transformed with the covariance matrix computed at each iteration (Figure 3), around 12% of the time is spent calculating the function studied. This number goes up to around 55% (4.4 times more) when using slice sampling real and computing the covariance function every $0.05K$ iterations (Figure 5). In both cases, the number of times the energy function was called is roughly the same (around 580000 times). This shows that, by performing slice sampling in the real space rather than in the transformed space and not computing the covariance matrix at every iteration, we removed calculations that are not needed and that bring a significant computational overhead: the computational time is reduced by a factor 2.8, as can be deduced from Table 2 which gives the computational time of all three cases for 20000 iterations.
@@
Figure 5: Profiling for slice sampling real with the calculation of the covariance matrix every $0.05K$ iterations
@@
For the impact of the parallelisation on the computation time, we also look at the time required to run nested_fit with and without parallelisation using slice sampling transformed with $K=1000$ and $n_{\textrm{bases}}=1$. The covariance matrix is computed every iteration. The parallelised case over 64 cores is around 21 times faster than the non-parallelised case (the non-parallelised case takes 85 minutes while the parallelised case takes 4 minutes). A possible explanation for this dividing factor being smaller than the number of cores used is that the time gained by simultaneously searching 64 new points may be partially counterbalanced by the time taken for the sequential adding of all the points, some of which may be rejected due to the constraint changing with each added point.
Table 1: Purpose of the functions shown in Figures 3 to 5.
@@
Performs one iteration of nested sampling with slice sampling real
@@
Version 4.5.5 - Slice Version 4.6.0 - Slice Version 4.6.0 - Slice
sampling transformed sampling transformed sampling real
@@ 5 Conclusion
First, we have studied clusters of two sizes. As for the 7-atom cluster, we have seen that we were able to recover phase transitions by comparing our work with that in Ref. partay_efficient_2010. We saw that we required more live points and more likelihood calls that in Ref. partay_efficient_2010, likely due to the different choices of method to find a new point, which is an inconvenience for a more computationally expensive energy function. We then studied the 36-atom cluster which presents a solid-solid phase transition at low temperature that nested_fit was able to recover provided enough live points are used.
We have then looked at the impact of different implementation on the computational resources: we have seen that using slice sampling transformed brought a consequent computational overhead through the need to transform back to the real space to compute the energy. Slice sampling real, which works directly in the real space, does not need to perform this task and therefore spends more than triple the time calculating the energy compared to slice sampling transformed, even though the function was called around the same number of times. We also saw that, using parallelisation on 64 cores, we gained a factor 21 in the computation time. All of those developments allowed us to study bigger and more complex clusters that require a higher number of live points. For example, one run for the 36-atom Lennard-Jones clusters with $K=70000$ took around 43 hours to run with the parallelisation.
@@
[…]
Conceptualisation, L.M., M.T. and F.F.; methodology, L.M., M.T. and F.F.; software, M.T., L.M. and C.G.; validation, L.M.; formal analysis, L.M.; investigation, L.M.; resources, F.F.; data curation, L.M. and M.T.; writing—original draft preparation, L.M.; writing—review and editing, F.F. and M.T..; visualisation, L.M.; supervision, M.T. and F.F.. All authors have read and agreed to the published version of the manuscript.
[…]
This research received no external funding.
[…]
Raw data can be provided on demand.
Acknowledgements.
[…]